# Randomization Tests for Conditional Group Symmetry

### v1

@@ Appendix B Proofs
    If $G\sim\lambda$ and $W\sim\rho_{*}P$ with $G{\;\perp\!\!\!\perp\;}W$, then $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GW$.
@@
which implies that ${X\mid\rho(X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)\mid\rho(X)}$.
@@
[…]
@@
Due to the sufficiency of $\rho(X)$ for $\mathcal{P}^{\circ}(\mathbf{X})$, the samples $X^{(b)}_{1:n}=(G^{(b)}_{1}X_{1},\dotsc,G^{(b)}_{n}X_{n})$ are conditionally i.i.d. (and therefore conditionally exchangeable) given $\rho(X)_{1:n}$, with the same conditional distribution as $X_{1:n}$ under $H_{0}$. Consequently, $\left(T(X^{(b)}_{1:n})\right)_{b=0}^{B}$ are conditionally i.i.d. given $\rho(X)_{1:n}$, with the same conditional distribution as $T(X_{1:n})$ under the null. It then follows from Lemma B.3 that (4) holds.
B.2 Proofs for Section 4
Proof of Theorem 4.1.
Define the random variable $\tilde{Y}\coloneqq\tilde{\tau}^{-1}Y$, where $\tilde{\tau}\sim P_{\tilde{\tau}|X}$, so that $P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}|\tilde{\tau},x)=\int\delta_{\tilde{\tau}^{-1}y}(d\tilde{y})P_{Y|X}(dy|x)$. For any integrable function $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$, by Fubini’s theorem,
$\int f(\tilde{\tau},x,\tilde{y})$ $P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f(\tilde{\tau},x,\tilde{\tau}^{-1}y)P_{Y|X}(dy\mid x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)\;.$
From this follows the identity $P_{\tilde{Y}|\tilde{\tau},X}(B|\tilde{\tau},x)=P_{Y|X}(\tilde{\tau}B|x)$.
Now assume that $P_{Y|X}$ is equivariant, so that for each $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{Y}}$, $P_{Y|X}(B|gx)=P_{Y|X}(g^{-1}B|x)$ for $P_{X}$-almost all $x$. Then for any $\tilde{\tau}\in\mathbf{G},x\in\mathbf{X},g\in\mathbf{G}$ and integrable $f\colon\mathbf{Y}\to\mathbb{R}$,
$\int P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},gx)f(\tilde{y})$ $=\int f((g\tilde{\tau})^{-1}y)P_{Y|X}(dy\mid gx)$
$=\int f(\tilde{\tau}^{-1}g^{-1}gy)P_{Y|X}(dy\mid x)$
$=\int f(\tilde{\tau}^{-1}y)P_{Y|X}(dy\mid x)$
$=\int f(\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)\;.$
This shows that the mapping $(\tilde{\tau},x)\mapsto P_{\tilde{Y}|\tilde{\tau},X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\tilde{\tau},x)$ is $\mathbf{G}$-invariant. Therefore, by Proposition A.1, for any measurable maximal invariant $\tilde{M}\colon\mathbf{G}\times\mathbf{X}\to\mathbf{M}$, there is a unique Markov probability kernel $\tilde{R}\colon\mathbf{M}\times\mathbf{S}_{\mathbf{Y}}\to[0,1]$ such that
$P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)=P_{\tilde{Y}|\tilde{M}}(B\mid\tilde{M}(\tilde{\tau},x))\;,\quad\tilde{\tau}\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$
@@ B.2 Proofs for Section 4
$P_{\tilde{Y}|\tilde{\tau},X}(B\mid\tilde{\tau},x)=\tilde{R}(B\mid M(x))\;,\quad\tilde{\tau}\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$ (23)
@@
$\int$ $f(\tilde{\tau},x,\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$=\int f(\tilde{\tau},x,\tilde{y})\tilde{R}(d\tilde{y}\mid M(x))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)\;.$
Conversely, assume that $(\tilde{\tau},X){\;\perp\!\!\!\perp\;}\tilde{\tau}^{-1}Y\mid M(X)$. Then (23) holds for $P_{X}\otimes P_{\tilde{\tau}|X}$-almost all $(\tilde{\tau},x)\in\mathbf{G}\times\mathbf{X}$. In particular, $P_{\tilde{Y}|\tilde{\tau},X}$ is $\mathbf{G}$-invariant in $(\tilde{\tau},x)$ for $P_{X}\otimes P_{\tilde{\tau}|X}$-almost all $(\tilde{\tau},x)$. Recall also that the inversion kernel $P_{\tilde{\tau}|X}$ is $\mathbf{G}$-equivariant. Therefore, for any integrable $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$ and any $g\in\mathbf{G}$,
$\int f(\tilde{\tau},x,y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}(\tilde{\tau}^{-1}y))P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},gx)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},g^{-1}x,\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid g\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid g^{-1}x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},g^{-1}x,g^{-1}\tilde{\tau}\tilde{y})P_{\tilde{Y}|\tilde{\tau},X}(d\tilde{y}\mid\tilde{\tau},x)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},g^{-1}x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\tilde{\tau},x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid gx)P_{Y|X}(dy\mid gx)P_{X}(dx)$
$\qquad=\int f(\tilde{\tau},x,g^{-1}y)P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(dy\mid gx)P_{X}(dx)\;.$
@@
$Q^{\prime}(B\mid x)\coloneqq\begin{cases}P_{Y|X}(B\mid x)&\text{if }x\notin N^{\times}\;,\\ \int_{\mathbf{G}}P_{\tilde{\tau}|X}(d\tilde{\tau}\mid x)P_{Y|X}(\tilde{\tau}^{-1}B\mid\tilde{\tau}^{-1}x)&\text{if }x\in N^{\times}\;.\end{cases}$
Since $P_{\tilde{\tau}|X}$ and $P_{Y|X}$ are probability kernels, so too is $Q^{\prime}$. It is also straightforward to show that $Q^{\prime}$ is $\mathbf{G}$-equivariant, so that $Q^{\prime}$ is a regular version of $P_{Y|X}$ that is $\mathbf{G}$-equivariant for all $x\in\mathbf{X}$, and equivalent to $P_{Y|X}$ up to the null set $N^{\times}$.
If there exists a measurable representative inversion (function) $\tau$, then the same proof holds with the inversion kernel $P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ substituted by $\delta_{\tau(x)}$, resulting in the simplified conditional independence statement in (12).
If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then $\tilde{Y}=Y$. Moreover, $\tilde{\tau}{\;\perp\!\!\!\perp\;}Y\mid X$ by construction, and therefore $(\tilde{\tau},X){\;\perp\!\!\!\perp\;}Y\mid M(X)$ is implied by $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. ∎
For the proof of the distributional identities in Proposition 4.2 that are equivalent to equivariance, the following lemma will be useful.
Lemma B.4.
Under the conditions of Proposition 4.2, $P_{Y|X}$ is $\mathbf{G}$-equivariant if and only if $(X,Y)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}(\Gamma\rho(X^{\prime}),\Gamma\tilde{Y})$, where $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}X^{\prime}$ and $\Gamma,\tilde{Y}$ are sampled as in (14).
@@
We will prove that the distributional identity implies that $P_{Y|X}$ is equivariant. The converse follows as a special case of the stronger statements in the proof of Proposition 4.2 below.
Fix arbitrary $g\in\mathbf{G}$. From (i) the definition of $P_{\Gamma|[X]}$, (ii) Fubini’s theorem, and (iii) the equivariance of the inversion kernel $P_{\tilde{\tau}|X}$, for any integrable $f\colon\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
@@
$=\int f(\gamma[x],\gamma\tilde{y})P_{\tilde{Y}|[X]}(d\tilde{y}\mid[x])P_{\Gamma|[X]}(d\gamma\mid[x])P_{[X]}(d[x])$
[…]

### v2

@@ Appendix B Further background on testing for distributional invariance
    If $G\sim\lambda$ and $R\sim\rho_{*}P$ with $G{\;\perp\!\!\!\perp\;}R$, then $X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}GR$.
@@
which implies that $X|\rho(X)\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}G\rho(X)|\rho(X)$}.
@@
$P(X\in A\mid\rho(X))=\Phi_{\rho(X)*}\lambda(A)\;,\quad A\in\mathbf{S}_{\mathbf{X}}\;,\quad P\text{-a.s.}$
@@
Proof of Theorem B.1.
Due to the sufficiency of $\rho(X)$ for $\mathcal{P}^{\circ}(\mathbf{X})$, the samples $X^{(b)}_{1:n}=(G^{(b)}_{1}X_{1},\dotsc,G^{(b)}_{n}X_{n})$ are conditionally i.i.d. (and therefore conditionally exchangeable) given $\rho(X)_{1:n}$, with the same conditional distribution as $X_{1:n}$ under $H_{0}$. Consequently, $\left(T(X^{(b)}_{1:n})\right)_{b=0}^{B}$ are conditionally i.i.d. given $\rho(X)_{1:n}$, with the same conditional distribution as $T(X_{1:n})$ under the null. It then follows from Lemma B.4 that (2) holds. ∎
Appendix C Proofs and additional theoretical results
C.1 Proof of Theorem 3.1
@@ Appendix C Proofs and additional theoretical results
Define the random variable $\tilde{Y}\coloneqq\Psi^{-1}Y$, where $\Psi|X\sim P_{\Psi|X}$, so that $P_{\tilde{Y}|\Psi,X}(d\tilde{y}|\psi,x)=\int\delta_{\psi^{-1}y}(d\tilde{y})P_{Y|X}(dy|x)$. For any integrable function $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$, by Fubini’s theorem,
[…]
$=\int f(\psi,x,\psi^{-1}y)P_{Y|X}(dy\mid x)P_{\Psi|X}(d\psi\mid x)P_{X}(dx)\;.$
From this follows the identity $P_{\tilde{Y}|\Psi,X}(B|\psi,x)=P_{Y|X}(\psi B|x)$, each $B\in\mathbf{S}_{\mathbf{Y}}$.
Now assume that $P_{Y|X}$ is equivariant, so that for each $g\in\mathbf{G},x\in\mathbf{X},B\in\mathbf{S}_{\mathbf{Y}}$, $P_{Y|X}(B|gx)=P_{Y|X}(g^{-1}B|x)$ for $P_{X}$-almost all $x$. Then for any $\psi\in\mathbf{G},x\in\mathbf{X},g\in\mathbf{G}$ and integrable $f\colon\mathbf{Y}\to\mathbb{R}$,
[…]
$=\int f(\psi^{-1}g^{-1}gy)P_{Y|X}(dy\mid x)$
$=\int f(\psi^{-1}y)P_{Y|X}(dy\mid x)$
$=\int f(\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid\psi,x)\;.$
This shows that the mapping $(\psi,x)\mapsto P_{\tilde{Y}|\Psi,X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|\psi,x)$ is $\mathbf{G}$-invariant. Therefore, by Proposition A.1, for any measurable maximal invariant $\tilde{M}\colon\mathbf{G}\times\mathbf{X}\to\mathbf{M}$, there is a unique Markov probability kernel $\tilde{R}\colon\mathbf{M}\times\mathbf{S}_{\mathbf{Y}}\to[0,1]$ such that
$P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid\psi,x)=P_{\tilde{Y}|\tilde{M}}(B\mid\tilde{M}(\psi,x))\;,\quad\psi\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$
@@
$P_{\tilde{Y}|\Psi,X}(B\mid\psi,x)=\tilde{R}(B\mid M(x))\;,\quad\psi\in\mathbf{G},\ x\in\mathbf{X},\ B\in\mathbf{S}_{\mathbf{Y}}\;.$ (18)
@@
$\int$ $f(\psi,x,\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid\psi,x)P_{\Psi|X}(d\psi\mid x)P_{X}(dx)$
$=\int f(\psi,x,\tilde{y})\tilde{R}(d\tilde{y}\mid M(x))P_{\Psi|X}(d\psi\mid x)P_{X}(dx)\;.$
Conversely, assume that $(\Psi,X){\;\perp\!\!\!\perp\;}\Psi^{-1}Y\mid M(X)$. Then (18) holds for $P_{X}\otimes P_{\Psi|X}$-almost all $(\psi,x)\in\mathbf{G}\times\mathbf{X}$. In particular, $P_{\tilde{Y}|\Psi,X}$ is $\mathbf{G}$-invariant in $(\psi,x)$ for $P_{X}\otimes P_{\Psi|X}$-almost all $(\psi,x)$. Recall also that the inversion kernel $P_{\Psi|X}$ is $\mathbf{G}$-equivariant. Therefore, for any integrable $f\colon\mathbf{G}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$ and any $g\in\mathbf{G}$,
$\int f(\psi,x,y)P_{\Psi|X}(d\psi\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\psi,x,\psi(\psi^{-1}y))P_{\Psi|X}(d\psi\mid x)P_{Y|X}(dy\mid x)P_{X}(dx)$
$\qquad=\int f(\psi,x,\psi\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid\psi,x)P_{\Psi|X}(d\psi\mid x)P_{X}(dx)$
$\qquad=\int f(\psi,x,\psi\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid g\psi,gx)P_{\Psi|X}(d\psi\mid x)P_{X}(dx)$
$\qquad=\int f(\psi,g^{-1}x,\psi\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid g\psi,x)P_{\Psi|X}(d\psi\mid g^{-1}x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\psi,g^{-1}x,g^{-1}\psi\tilde{y})P_{\tilde{Y}|\Psi,X}(d\tilde{y}\mid\psi,x)P_{\Psi|X}(d\psi\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\psi,g^{-1}x,g^{-1}y)P_{\Psi|X}(d\psi\mid x)P_{Y|X}(dy\mid x)(g_{*}P_{X})(dx)$
$\qquad=\int f(g^{-1}\psi,x,g^{-1}y)P_{\Psi|X}(d\psi\mid gx)P_{Y|X}(dy\mid gx)P_{X}(dx)$
$\qquad=\int f(\psi,x,g^{-1}y)P_{\Psi|X}(d\psi\mid x)P_{Y|X}(dy\mid gx)P_{X}(dx)\;.$
@@
$Q^{\prime}(B\mid x)\coloneqq\begin{cases}P_{Y|X}(B\mid x)&\text{if }x\notin N^{\times}\;,\\ \int_{\mathbf{G}}P_{\Psi|X}(d\psi\mid x)P_{Y|X}(\psi^{-1}B\mid\psi^{-1}x)&\text{if }x\in N^{\times}\;.\end{cases}$
Since $P_{\Psi|X}$ and $P_{Y|X}$ are probability kernels, so too is $Q^{\prime}$. It is also straightforward to show that $Q^{\prime}$ is $\mathbf{G}$-equivariant, so that $Q^{\prime}$ is a regular version of $P_{Y|X}$ that is $\mathbf{G}$-equivariant for all $x\in\mathbf{X}$, and equivalent to $P_{Y|X}$ up to the null set $N^{\times}$.
If there exists a measurable representative inversion (function) ${\psi^{\circ}}({\,\vbox{\hbox{\tiny$\bullet$}}\,})$, then the same proof holds with the inversion kernel $P_{\Psi|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|x)$ substituted by $\delta_{{\psi^{\circ}}(x)}$, resulting in the simplified conditional independence statement in (5).
If the action of $\mathbf{G}$ on $\mathbf{Y}$ is trivial, then $\tilde{Y}=Y$. Moreover, $\Psi{\;\perp\!\!\!\perp\;}Y\mid X$ by construction, and therefore $(\Psi,X){\;\perp\!\!\!\perp\;}Y\mid M(X)$ is implied by $X{\;\perp\!\!\!\perp\;}Y\mid M(X)$. ∎
C.2 Further characterizations of conditional equivariance
Theorem 3.1 can be refined to obtain distributional identities that underpin the randomization tests of Section 4. Proofs follow Corollary C.3 below.
Proposition C.1.
Let $P_{X,Y}=P_{X}\otimes P_{Y|X}$ be a probability measure on $\mathbf{X}\times\mathbf{Y}$. For a lcscH group $\mathbf{G}$ acting on $\mathbf{X}$ and $\mathbf{Y}$, with the action on $\mathbf{X}$ proper, let $P_{\Gamma|R}$ be the group-orbit kernel defined in (8). Then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E1.
[…]
  2. E2.
[…]
  3. E3.
[…]
Proposition C.1 has two special cases that are important for practical reasons. First, if $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$, the data being used in the test can be reduced to $Y$ only.
Corollary C.2.
Assume the conditions of Proposition C.1. If $\mathbf{G}$ acts non-trivially on $\mathbf{Y}$ then the following statements are equivalent to $P_{Y|X}$ being $\mathbf{G}$-equivariant.
  1. E4.
[…]
  2. E5.
[…]
  3. E6.
[…]
Corollary C.3.
  1. E7.
[…]
  2. E8.
For the proof of the distributional identities in Proposition C.1 that are equivalent to equivariance, the following lemma will be useful.
Lemma C.4.
[…]
@@ C.2 Further characterizations of conditional equivariance
We will prove that the distributional identity implies that $P_{Y|X}$ is equivariant. The converse follows as a special case of the stronger statements in the proof of Proposition C.1 below.
Fix arbitrary $g\in\mathbf{G}$. From (i) the definition of $P_{\Gamma|R}$, (ii) Fubini’s theorem, and (iii) the equivariance of the inversion kernel $P_{\Psi|X}$, for any integrable $f\colon\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
@@
$=\int f(\gamma r,\gamma\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{\Gamma|R}(d\gamma\mid r)P_{R}(dr)$
$\overset{\text{(i),(ii)}}{=}\int f(\psi r,\psi\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{\Psi|X}(d\psi\mid x^{\prime})P_{X|R}(dx^{\prime}\mid r)P_{R}(dr)$
$\overset{\text{(iii)}}{=}\int f(g\psi r,g\psi\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{\Psi|X}(d\psi\mid g^{-1}x^{\prime})P_{X|R}(dx^{\prime}\mid r)P_{R}(dr)$
@@
Proof of Proposition C.1.
Let $(X,Y)\sim P_{X,Y}$. Assume that $P_{Y|X}$ is
$\mathbf{G}$-equivariant. To show Property E1, for any integrable $f\colon\mathbf{R}\times\mathbf{Y}\times\mathbf{X}\times\mathbf{Y}\to\mathbb{R}$,
$\int$ $f(\rho(x),\tilde{y},x,y)P(dx,d\tilde{y},dy)$
$=\int f(r,\tilde{y},x^{\prime},y)P_{Y|X}(dy\mid\tilde{y},x^{\prime})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{X|R}(dx^{\prime}\mid r)P_{R}(dr)$
$=\int f(r,\tilde{y},\psi r,\psi\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{\Psi|X}(d\psi\mid x^{\prime})P_{X|R}(dx^{\prime}\mid r)P_{R}(dr)$
$=\int f(r,\tilde{y},\gamma r,\gamma\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid r)P_{\Gamma|R}(d\gamma\mid r)P_{R}(dr)$
$=\int f(\rho(x),\tilde{y},\gamma\rho(x),\gamma\tilde{y})P_{\tilde{Y}|R}(d\tilde{y}\mid\rho(x))P_{\Gamma|R}(d\gamma\mid\rho(x))P_{X}(dx)\;,$
[…]
Proof of Corollary C.2.
Property E4Property E6 follow as special cases of the corresponding properties in Proposition C.1, marginalizing over $X$. ∎
Proof of Corollary C.3.
Properties E7 and E8 follow from the fact that
$P_{\Gamma|R}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|r)=\lambda({\,\vbox{\hbox{\tiny$\bullet$}}\,})$ for each $r$ if and only if $P_{X}$ is $\mathbf{G}$-invariant. ∎
C.3 Proofs for Section 4